we burn 345 calories hiking in 45 min. preparing for a 6 hr hike with special beverage it gives 310 calories replace 1/3 of the calories burned carrying the least packs how many packs do i take

Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(52832)

.

I understood not much from reading this text - so inaccurately it is written.

But I found the same problem in the Internet, written clearly and accurately under this link
https://brainly.com/question/17186310

Here is the text from that web-page

    Jayden, who burns 345 calories in 45 min while hiking is preparing for a 6 hour hike.
    He uses a special supplement beverage pack that provides water, needed electrolytes, and 310 calories. 
    The goal is to replace roughly 1/3 of the calories burned while carrying as light a load as possible.
    How many packs should he take?

Below is the solution from there, in one line (of 08/11/2020)
    I think that he should bring three packs. 
    6 hours = 360 minutes. 360/45 = 8. 345x8 = 2760. 2760/3 = 920 and 310x3 = 930. :)


I am surprised how it is possible to write so sloppily as the visitor does in his post.



Answer by Theo(13342)

you burn 345 calories hiking for 45 minutes.
assuming the calorie burn rate is the same throughout any longer hike, then you are turning calories at the rate of 345/45 calories per minute.

if you are hiking for 6 hours, then you will burn 345/45 calories per minute for 6 * 60 = 360 minutes.
the total number of calories consumed would be 345/45 per minute for 360 minutes = 2760 calories.

if the special beverage gives you 310 calories per pack, and you want to replace all of the calories consumed, then you would need to take 2760 / 310 = 9 packs of beverage.

if you only want to replace 1/3 of the calories consumed, then you would need to take (2760 / 3) / 310 = 3 packs of beverage.

9 packs of beverage would give you 310 * 9 = 2790 calories, which is more than the 2760 consumed.

3 packs of beverage would give you 310 * 3 = 930 calories, which is more than the 2760/3 = 920 calories consumed.

let me know if you have any questions.

theo